# A primer on quasi-convex functions in nonlinear potential theories

### v2

@@ 2.4. The theorems on upper contact jets and on summands
For the sake of completeness, in Subsection 4.3, we will give a proof of the classical Theorem on Sums for semicontinuous functions (see Theorem 4.12). It is easily deduced from its quasi-convex counterpart because, roughly speaking, it is possible to transform a function which is bounded above into a quasi-convex function in such a way that it produces a regular displacement of the contact points of every given upper contact jet. More precisely, the displacement is proportional to the gradient-component of the jet, as we will see in Subsection 4.2. In addition, in Subsection 9.2 we will introduce and prove a potential theoretic version of Harvey and Lawson in Theorem 9.20, which they used in their proof of strict comparison principles with minimal monotonicity for constant coefficient potential theories.
3. The lemmas of Jensen and Słodkowski
@@
In the crucial passage “from almost everywhere to everywhere”, two classical instruments dealing with (quasi-)convex functions are Słodkowski’s density estimate [43, Theorem 3.2] and Jensen’s Lemma [35, Lemma 3.10]. The first instrument yields Słodkowski’s Largest Eigenvalue Theorem [43, Corollary 3.5] which, in turn, plays an essential role in Harvey and Lawson’s proof of the Subaffine Theorem [23, Theorem 6.5]. The second instrument yields Jensen’s maximum principle [35, Theorem 3.1] which, in turn, is a key ingredient in the proof of the Theorem on Sums [13, Theorem 3.2]. These two instruments were developed independently. The first in the context of pluripotential theory and the second in the context of machinery for proving the uniqueness of viscosity solutions. As we will see, they are in some sense equivalent as was shown by Harvey and Lawson. This long unrecognized equivalence is a good example of a missed opportunity for synergy between potential theory and operator theory where such a synergy is a main theme of the program initiated by Harvey and Lawson.
@@
Słodkowski’s original proof of his density estimate is based on two nontrivial properties [43, Proposition 3.3(iii) and Lemma 3.4]. The first property concerns the largest generalized eigenvalue of the Hessian at $x$ of a convex function $u$ [43, Definition 3.1] and the second property concerns spheres which support from above the graph $\Gamma(u)$ of $u$ at the point $(x,u(x))$; that is, spheres in ${\mathbb{R}}^{n+1}$ lying above $\Gamma(u)$ and touching it only at $(x,u(x))$. Following Harvey and Lawson [25], our proof of Słodkowski’s estimate will be based on the simpler “paraboloidal” counterparts of such properties.
@@
At the end of the section, a second non classical proof of Jensen’s lemma will be proposed. Jensen points out that his lemma is based ideas of Pucci [39] and gives a proof which relies on the area formula (see, e.g., [19, Theorem 3.2.3]); we will offer a different area-formula-based proof by reformulating Jensen’s lemma à la Harvey–Lawson [25] and then following an argument by Harvey [20] which passes also through Alexandrov’s maximum principle (see, for example, [7]).
3.1. Classical statements: a review
For the convenience of the reader, we recall some classical statements of the Jensen and Słodkowki lemmas which we deal with.
@@ 3.1. Classical statements: a review
if $x\in\operatorname{Diff}^{1}\!u$ and $\mathcal{K}(u,x)\vcentcolon=+\infty$ otherwise, and proves the following estimate (see [43, Theorem 3.2]). We will give an original “paraboloidal” proof of this result as a corollary of our novel Lemma 3.21.
@@
Remark 3.2.
$\mathcal{K}(u,x)$ indeed generalizes the concept of largest eigenvalue of the Hessian of a convex function $u$ at $x$, since, if $u$ has second-order Peano derivatives at $x$, then $\mathcal{K}(u,x)=\lambda_{n}(D^{2}u(x))$, the largest eigenvalue of the Hessian $D^{2}u(x)$:
The two nontrivial properties on which Słodkowski’s proof of the density estimate are recalled in the following two lemmas which correspond to Proposition 3.3(iii) and Lemma 3.4, resepectively, in [43].
@@
Let $U\subset{\mathbb{R}}^{n}$ be open and $u\colon U\to{\mathbb{R}}$ be convex. Suppose there exists a sphere $\mathbb{S}(c,r)\vcentcolon=\partial B_{r}(c)\subset{\mathbb{R}}^{n+1}$ which supports $\Gamma(u)$ from the above at $(x,u(x))$, with $x\in\operatorname{Diff}^{1}\!u$; then
@@
We will formulate and prove an original “paraboloidal” version of this upper bound in Lemma 3.21, which exploits the paraboloidal machinery developed by Harvey and Lawson in [25] and leads to the aforementioned paraboloidal proof of Lemma 3.1.
@@
Let $u$ be nonnegative and convex on $B_{\rho}\subset{\mathbb{R}}^{n}$, for some $\rho>0$, with $u(0)=0$ and $Du(0)=0$. Assume that there exists a ball $B_{R}((0,\dots,0,R))\subset{\mathbb{R}}^{n+1}$ which intersects the graph of $u$ only at $0\in{\mathbb{R}}^{n+1}$, and let, for $0<r<R$,
@@
As a corollary of his density estimate in Lemma 3.1, Słodkowski obtains the following interesting result (see Corollary 3.5 of [43]).
@@
Let $u$ be locally convex on an open set $X$ and suppose $K(u,x)\geqslant M$ for a.e. $x\in X$. Then $K(u,x)\geqslant M$ for all $x\in X$.
We will give a proof of this result under the weaker hypothesis that $u$ is just locally quasi-convex in Theorem 3.30 as a consequence of the Jensen-Słodkowski Theorem.
We conclude this subsection by recalling Jensen’s lemma, which we present as one finds in Lemma A.3 of the User’s guide [13].
@@
Let $w\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ be quasi-convex and $\hat{x}$ be a strict local maximum point of $w$. For $p\in{\mathbb{R}}^{n}$, set $w_{p}(x)=w(x)+\langle p,x\rangle$. Then, for $r,\delta>0$, the set
@@
Jensen’s original statement [35, Lemma 3.10] uses seemingly different hypotheses; that is, Jensen assumes that $u$ is continuous, $u$ belongs to $W^{1,\infty}$ and satisfies $D_{\nu}^{2}u\geqslant-\lambda I$ in the sense of distributions, for all directions $\nu$. Nevertheless, one can see that these requirements are equivalent to asking that $u$ be $\lambda$-quasi-convex (for instance, one can use [16, Theorem 31]).
@@ 3.2. Upper contact quadratic functions and the vertex map
Our journey towards the reformulation of the classical results of Jensen and Słodkowski begins with a presentation the “paraboloidal” machinery of of Harvey-Lawson [25] which substitutes the original “spherical” machinery of Słodkowski [43]. This machinery concerns upper quadratic test functions of a given radius, upper contact points and jets of type $A$ (with fixed matrix component $A\in\mathcal{S}(n)$) and the important associated upper vertex map. In the next subsection, these notions will lead to estimates on the measure of contact points of fixed type.
@@
We say that $\varphi$ is a quadratic function of radius $r$ if $D^{2}\varphi\equiv\frac{1}{r}I$. In this case it is easy to see that $\varphi$ can be written in a unique way as $\varphi=\varphi(v)+\frac{1}{2r}|\cdot-v|^{2}$; we will call $v$ the vertex point of $\varphi$.
@@
We recall that in Definition 2.6, a point $x\in X$ for which (3.2) holds for some quadratic function $\varphi$ was called an upper contact point for $u$ and we presented the upper contact quadratic functions in the form $\varphi(x)=u(x)+\langle p,y-x\rangle+Q_{A}(y-x)$ where $Q_{A}(x)=\frac{1}{2}\langle Ax,x\rangle$ and the contact condition (3.2) takes the form (2.4); that is,
@@
Note that $\varphi$ is the (unique) quadratic function with $J^{2}_{x}(x)=(u(x),p,A)$ and $(p,A)\in J^{2,+}_{x}u$ was called an upper contact jet for $u$ in $x$. Two variants of upper contact jets will be used; namely those with fixed matrix component $A\in\mathcal{S}(n)$ and then also those for which (3.3) holds on a fixed subset of $X$.
@@
Remark 3.11 (Notation).
We will denote by $B_{\rho}(x)\subset{\mathbb{R}}^{n}$ the open ball of radius $\rho$ about $x$, and we will omit the center when it is the origin; that is, we will use the notation $B_{\rho}\vcentcolon=B_{\rho}(0)$.
@@
We are going to focus our attention on sets of upper contact points of fixed type which are global on closed balls; that is of the form $\mathcal{C}^{+}(u,\overline{B}_{\rho},A)$, up to translations that center the ball at the origin. In particular, we will be mainly interested in the Lebesgue measure of such sets, so let us also point out that, when it comes to measures, if $u$ is continuous, then one can equivalently consider open or closed balls, since
@@
We are going to discuss the Lebesgue measure of the set $\mathcal{C}^{+}(u,X,A)$ in the special case when $u$ is locally quasi-convex. For the sake of completeness, we note that $\mathcal{C}^{+}(u,X,A)$ is a measurable set when $u$ is locally quasi-convex. Indeed, if one defines $h_{y,p,A}\vcentcolon=u-u(y)+\langle p,y-\,\cdot\rangle+Q_{A}(y-\,\cdot)$, then it is easy to see that
@@
The following result, which is Lemma 5.1 of [25], is an interesting exercise that shows an elementary geometric property of the convex hull of two open paraboloids (namely the open epigraphs of two quadratic functions). It will play a key role in Harvey and Lawson’s paraboloidal treatment of the Słodkowski Lemma 3.1. We will denote the graph of $\varphi$ by $\Gamma(\varphi)$, its open epigraph by $\operatorname{epi}(\varphi)$, and the convex hull of a subset $S\subset{\mathbb{R}}^{n+1}$ by $\operatorname{conv}(S)$.
@@
There is an open vertical slab $\mathfrak{S}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}$ written as the intersection $\mathcal{H}_{1}\cap\mathcal{H}_{2}$ of two parallel vertical open half-spaces with the following property. Let :$\mathfrak{C}\vcentcolon=\operatorname{conv}(\operatorname{epi}(\varphi_{1})\cup\operatorname{epi}(\varphi_{2}))$, then
[…]
@@
Let $z\in\partial\big{(}\!\operatorname{epi}(\varphi_{1})\cup\operatorname{epi}(\varphi_{2})\big{)}\setminus\big{(}\Gamma(\varphi_{1})\cap\Gamma(\varphi_{2})\big{)}=\vcentcolon\mathfrak{B}$ and let $H_{z}$ be the tangent hyperplane to $\operatorname{epi}(\varphi_{1})\cup\operatorname{epi}(\varphi_{2})$ at $z$. Then one of the following is verified (see Figure 1):
@@
    $H_{z}\cap\partial(\operatorname{epi}(\varphi_{1})\cup\operatorname{epi}(\varphi_{2}))$ is a singleton;
@@
    $H_{z}\cap\partial(\operatorname{epi}(\varphi_{1})\cup\operatorname{epi}(\varphi_{2}))$ is a doubleton;
@@
    $H_{z}\cap\partial(\operatorname{epi}(\varphi_{1})\cup\operatorname{epi}(\varphi_{2}))$ is infinite.
@@
$\partial\mathfrak{S}\cap\partial\!\bigcup_{j=1,2}\!\operatorname{epi}(\varphi_{j})=\{z:H_{z}\ \text{satisfies \emph{(ii)}}\}=\bigcup_{j=1,2}(\partial\mathcal{H}_{j}\cap\Gamma(\varphi_{j})),$
@@
Figure 1. The three situations described in Remark 3.15.
@@
Figure 2. The slab constructed in the proof of Lemma 3.14.
The slabbed hull lemma will be used to prove a crucial property of the vertex map, which is the main tool we will need. In preparation for the definition of this map, we note that a generic upper contact quadratic function of radius $r$ for $u$ at $x$ takes the form $\varphi(\cdot)=u(x)+\langle p,\cdot-x\rangle+\frac{1}{2r}|\cdot-\,x|^{2}$ and has vertex $v=x-rp$. Moreover, if $x\in\mathrm{Diff}^{1}(u)$, then $p=Du(x)$ is unique. This motivates the following definition.
@@
$V_{r}\colon\mathcal{C}^{+}(u,X,\tfrac{1}{r}I)\cap\mathrm{Diff}^{1}(u)\to{\mathbb{R}}^{n},\quad\text{with}\ \ V_{r}\vcentcolon=I-rDu,$
@@
Note that the domain of $V_{r}$ consists of those points $x\in X$ which admit a global upper contact quadratic function of radius $r$ and for which $u$ is differentiable. If $u$ is (locally quasi-)convex, then the upper vertex map is well-defined on all of $\mathcal{C}^{+}(u,X,\tfrac{1}{r}I)$ since $\mathcal{C}^{+}(u,X,\frac{1}{r}I)\subset\mathrm{Diff}^{1}(u)$ by Lemma 2.13 (locally quasi-convex functions are differentiable at all upper contact points).
@@
For $j=1,2$, given $x_{j}\in\mathcal{C}^{+}(u,X,\frac{1}{r}I)$, let $\varphi_{j}$ be the upper contact quadratic function of radius $r$ for $u$ at $x_{j}$ with vertex $v_{j}\vcentcolon=V_{r}(x_{j})$. With the notation in Lemma 3.14, since $u$ is convex and whose graph lies below those of the $\varphi_{j}$’s, then $\mathfrak{C}\subset\operatorname{epi}(u)$. Recall that $\operatorname{epi}$ is the open epigraph, so that $z_{j}\vcentcolon=(x_{j},\varphi_{j}(x_{j}))=(x_{j},u(x_{j}))\in\Gamma(\varphi_{j})\setminus\operatorname{epi}(u)$; hence $z_{j}\in\Gamma(\varphi_{j})\setminus\mathfrak{C}$. Therefore, by Lemma 3.14, $z_{j}\notin\mathcal{H}_{j}$. It follows that the $z_{j}$’s lie on opposite sides of $\mathfrak{S}$, therefore $|x_{1}-x_{2}|\geqslant|v_{1}-v_{2}|$. ∎
@@
[…]
@@
is nonempty and compact.
Since $u$ is bounded below, its oscillation is finite. Suppose that $\delta$ is small enought to ensure that $K_{\delta}$ is not empty and fix $v\in K_{\delta}$. Clearly, if $x\in\mathcal{C}^{+}(u,K,\frac{1}{r}I,v)$, then

### v3

@@ 2.4. The theorems on upper contact jets and on summands
For the sake of completeness, in Section 4.3, we will give a proof of the classical Theorem on Sums for semicontinuous functions (see Theorem 4.13). It is easily deduced from its semiconvex counterpart because, roughly speaking, it is possible to transform a function which is bounded above into a semiconvex function in such a way that it produces a regular displacement of the contact points of every given upper contact jet. More precisely, the displacement is proportional to the gradient-component of the jet, as we will see in Section 4.2. In addition, in Section 9.3 we will introduce and prove a potential theoretic version of Harvey and Lawson in Theorem 9.22, which they used in their proof of strict comparison principles with minimal monotonicity for constant coefficient potential theories.
Chapter 3 The lemmas of Jensen and Słodkowski
In this chapter we treat the remaining deep analytical results which are needed for a robust viscosity theory for subharmonics in general potential theories and subsolutions of fully nonlinear ellipitc PDEs. We address the central problem of how to ensure a priori the existence of a sufficient amount of upper contact points for locally semiconvex functions which, when combined with Alexandrov’s Theorem and the sup-convolution approximation of semicontinuous functions by semiconvex functions, will complete the foundations of the “hard analysis” part of the viscosity theory. The analysis will culminate in the so-called Jensen–Słodkowski Theorem 3.27, which unifies and generalizes two historically distinct approaches (Jensen’s Lemma and Słodkowski’s density estimate) to the central problem in a viscosity approach to potential theory and operator theory. Several important consequences of the Jensen–Słodkowski Theorem are also given, as well as a novel proof of this fundamental theorem which is shown to follow from the area formula for gradients of locally semiconvex functions.
3.1. Classical statements: a review
@@
In the crucial passage “from almost everywhere to everywhere” in the viscosity theory, two classical instruments dealing with (semi-)convex functions are Słodkowski’s density estimate [48, Theorem 3.2] and Jensen’s Lemma [39, Lemma 3.10]. The first instrument yields Słodkowski’s Largest Eigenvalue Theorem [48, Corollary 3.5] which, in turn, plays an essential role in Harvey and Lawson’s proof of the Subaffine Theorem [26, Theorem 6.5]. The second instrument yields Jensen’s maximum principle [39, Theorem 3.1] which, in turn, is a key ingredient in the proof of the Theorem on Sums [14, Theorem 3.2]. These two instruments were developed independently. The first in the context of pluripotential theory and the second in the context of the machinery for proving the uniqueness of viscosity solutions. As we will see, they are in some sense equivalent. This long unrecognized equivalence is a good example of a missed opportunity for synergy between potential theory and operator theory where such a synergy is a main theme of the program initiated by Harvey and Lawson.
@@
Słodkowski’s original proof of his density estimate is based on two nontrivial properties [48, Proposition 3.3(iii) and Lemma 3.4]. The first property concerns the largest generalized eigenvalue of the Hessian at $x$ of a convex function $u$ [48, Definition 3.1] and the second property concerns spheres which support from above the graph $\Gamma(u)$ of $u$ at the point $(x,u(x))$; that is, spheres in ${\mathbb{R}}^{n+1}$ lying above $\Gamma(u)$ and touching it only at $(x,u(x))$. Following Harvey and Lawson [28], our proof of Słodkowski’s estimate will be based on the simpler “paraboloidal” counterparts of such properties.
@@
At the end of the section, a second non classical proof of Jensen’s lemma will be proposed. Jensen points out that his lemma is based ideas of Pucci [44] and gives a proof which relies on the area formula (see, e.g., [20, Theorem 3.2.3]); we will offer a different area-formula-based proof by reformulating Jensen’s lemma à la Harvey–Lawson [28] and then following an argument by Harvey [23] which passes also through Alexandrov’s maximum principle (see, for example, [7]).
For the convenience of the reader, we recall some classical statements of Jensen’s and Słodkowski’s lemmas which we deal with.
@@
if $x\in\operatorname{Diff}^{1}\!u$ and $\mathcal{K}(u,x)\vcentcolon=+\infty$ otherwise.
Remark 3.1.
$\mathcal{K}(u,x)$ indeed generalizes the concept of largest eigenvalue of the Hessian of a convex function $u$ at $x$, since, if $u$ has second-order Peano derivatives at $x$, then $\mathcal{K}(u,x)$ is the largest eigenvalue of the Hessian $D^{2}u(x)$:
Słodkowski’s density estimate is stated next. See [48, Theorem 3.2] for the original proof. We will give a novel “paraboloidal” proof of this result, which will be a corollary of our Lemma 3.20.
@@
The two nontrivial properties on which Słodkowski’s proof of the density estimate are recalled in the following two lemmas which correspond, respectively, to Proposition 3.3(iii) and Lemma 3.4 in [48].
@@
Let $u\colon U\to{\mathbb{R}}$ be convex on $U\subset{\mathbb{R}}^{n}$ open and convex. Suppose there exists a sphere $\mathbb{S}(c,r)\vcentcolon=\partial B_{r}(c)\subset{\mathbb{R}}^{n+1}$ which supports the graph of $u$ from the above at $(x,u(x))$ with $x\in\operatorname{Diff}^{1}\!u$. Then one has
@@
We will formulate and prove an original paraboloidal version of this upper bound in Lemma 3.20, which exploits the paraboloidal machinery developed by Harvey and Lawson in [28] and leads to the aforementioned paraboloidal proof of Lemma 3.2.
@@
Let $u$ be nonnegative and convex on $B_{\rho}\subset{\mathbb{R}}^{n}$, for some $\rho>0$, with $u(0)=0$ and $Du(0)=0$. Assume that there exists a ball $B_{R}((0,\dots,0,R))\subset{\mathbb{R}}^{n+1}$ which intersects $\Gamma(u)$ (the graph of $u$) only at $0\in{\mathbb{R}}^{n+1}$, and let, for $0<r<R$,
@@
As a corollary of his density estimate in Lemma 3.2, Słodkowski obtains the following interesting result (see [48, Corollary 3.5]).
@@
Let $u$ be locally convex on an open set $X$ and suppose $\mathcal{K}(u,x)\geqslant M$ for a.e. $x\in X$. Then $\mathcal{K}(u,x)\geqslant M$ for all $x\in X$.
We will give a proof of this result under the weaker hypothesis that $u$ is just locally semiconvex in Theorem 3.29 as a consequence of the Jensen–Słodkowski Theorem 3.27.
We conclude this section by recalling Jensen’s lemma, which we present as one finds in Lemma A.3 of the User’s guide [14].
@@
Let $w\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ be semiconvex and $\hat{x}$ be a strict local maximum point of $w$. For $p\in{\mathbb{R}}^{n}$, set $w_{p}(x)=w(x)+\langle p,x\rangle$. Then, for $r,\delta>0$, the set
@@
Jensen’s original statement [39, Lemma 3.10] uses seemingly different hypotheses; that is, Jensen assumes that $u$ is continuous, $u$ belongs to $W^{1,\infty}$ and satisfies $D_{\nu}^{2}u\geqslant-\lambda I$ in the sense of distributions, for all directions $\nu$. Nevertheless, one can see that these requirements are equivalent to asking that $u$ be $\lambda$-semiconvex (for instance, one can use [16, Theorem 3.1]).
@@ 3.2. Upper contact quadratic functions and the vertex map
Our journey towards the reformulation of the classical results of Jensen and Słodkowski begins with a presentation the “paraboloidal” machinery of Harvey–Lawson [28] which substitutes the original “spherical” one of Słodkowski [48]. This machinery concerns upper quadratic test functions of a given radius, upper contact points and jets of type $A$ (meaning with fixed matrix component $A\in\mathcal{S}(n)$) and the important associated upper vertex map. In the next section, these notions will lead to estimates on the measure of contact points of fixed type.
@@
We say that a function $\varphi$ is a quadratic function of radius $r$ if it is quadratic and satisfies $D^{2}\varphi\equiv\frac{1}{r}I$. In this case it is easy to see that $\varphi$ can be written in a unique way as $\varphi=\varphi(v)+\frac{1}{2r}|\cdot-v|^{2}$; we will call $v$ the vertex point of $\varphi$.
@@
We recall that in Definition 2.6, a point $x\in X$ for which (3.2) holds for some quadratic function $\varphi$ was called an upper contact point for $u$ and we presented the upper contact quadratic functions in the form $\varphi(x)=u(x)+\langle p,y-x\rangle+Q_{A}(y-x)$ where $Q_{A}(x)=\frac{1}{2}\langle Ax,x\rangle$ and the contact condition (3.2) takes the form
@@
Note that $\varphi$ is the (unique) quadratic function with $J^{2}_{x}\varphi=(u(x),p,A)$, and $(p,A)\in J^{2,+}_{x}u$ was called an upper contact jet for $u$ in $x$.
Two variants of upper contact jets will be used; namely those with fixed matrix component $A\in\mathcal{S}(n)$ and then also those for which (3.3) holds on a fixed subset of $X$.
@@
We are going to focus our attention on sets of upper contact points of fixed type which are global on closed balls centered at the origin; that is, of the form $\mathcal{C}^{+}(u,\overline{B}_{\rho},A)$, up to translations that center the ball at the origin. In particular, we will be mainly interested in the Lebesgue measure of such sets, so let us also point out that, when it comes to measures, if $u$ is continuous, then one can equivalently consider open or closed balls, since
@@
We are going to discuss the Lebesgue measure of the set $\mathcal{C}^{+}(u,X,A)$ in the special case when $u$ is locally semiconvex. For the sake of completeness, we note that $\mathcal{C}^{+}(u,X,A)$ is a measurable set when $u$ is locally semiconvex. Indeed, if one defines $h_{y,p,A}\vcentcolon=u-u(y)+\langle p,y-\,\cdot\rangle+Q_{A}(y-\,\cdot)$, then it is easy to see that
@@
The following result, which is to be found in [28, Lemma 5.1], is an interesting exercise that shows an elementary geometric property of the convex hull of two open paraboloids (namely the strict epigraphs of two quadratic functions). It will play a key role in Harvey and Lawson’s paraboloidal treatment of Słodkowski’s Lemma 3.2. As before, we will denote the graph of $\varphi$ by $\Gamma(\varphi)$, indicate its strict epigraph by $\operatorname{epi}_{S}(\varphi)$, and the convex hull of a subset $E\subset{\mathbb{R}}^{n+1}$ by $\operatorname{conv}(E)$. Recall that these objects are defined as follows: $\Gamma(\varphi)\vcentcolon=\{(x,y)\in{\mathbb{R}}^{n+1}:\,y=\varphi(x)\}$, $\operatorname{epi}_{S}(\varphi)\vcentcolon=\operatorname{epi}(\varphi)\setminus\Gamma(\varphi)$, and $\operatorname{conv}(E)\vcentcolon=\bigcap_{E\subset C\,\text{convex}}C$.
@@
Let :$\mathfrak{C}\vcentcolon=\operatorname{conv}(\operatorname{epi}_{S}(\varphi_{1})\cup\operatorname{epi}_{S}(\varphi_{2}))$. There is an open vertical slab $\mathfrak{S}\subset{\mathbb{R}}^{n}\times{\mathbb{R}}$, written as the intersection $\mathcal{H}_{1}\cap\mathcal{H}_{2}$ of two parallel vertical open half-spaces $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$, with the following property:
$\Gamma(\varphi_{j})\cap\mathfrak{C}=\Gamma(\varphi_{j})\cap\mathcal{H}_{j}\qquad\text{for}\ j=1,2.$
@@
Let
$z\in\partial\big{(}\!\operatorname{epi}_{S}(\varphi_{1})\cup\operatorname{epi}_{S}(\varphi_{2})\big{)}\setminus\big{(}\Gamma(\varphi_{1})\cap\Gamma(\varphi_{2})\big{)}=\vcentcolon\mathfrak{B}$
and let $H_{z}$ be the tangent hyperplane to $\operatorname{epi}_{S}(\varphi_{1})\cup\operatorname{epi}_{S}(\varphi_{2})$ at $z$. Then one of the following is verified (see Figure 3.1):
@@
    $H_{z}\cap\partial(\operatorname{epi}_{S}(\varphi_{1})\cup\operatorname{epi}_{S}(\varphi_{2}))$ is a singleton;
@@
    $H_{z}\cap\partial(\operatorname{epi}_{S}(\varphi_{1})\cup\operatorname{epi}_{S}(\varphi_{2}))$ is a doubleton;
@@
    $H_{z}\cap\partial(\operatorname{epi}_{S}(\varphi_{1})\cup\operatorname{epi}_{S}(\varphi_{2}))$ is infinite.
@@
$\partial\mathfrak{S}\cap\partial\!\bigcup_{j=1,2}\!\operatorname{epi}_{S}(\varphi_{j})=\{z:H_{z}\ \text{satisfies \emph{(ii)}}\}=\bigcup_{j=1,2}(\partial\mathcal{H}_{j}\cap\Gamma(\varphi_{j})),$
@@
Figure 3.1. The three situations described in Remark 3.14.
@@
Figure 3.2. The slab constructed in the proof of Lemma 3.13.
The slabbed hull property of Lemma 3.13 will be used to prove a crucial property of the vertex map, which is the main tool we will need. In preparation for the definition of this map, we note that a generic upper contact quadratic function of radius $r$ for $u$ at $x$ takes the form $\varphi(\cdot)=u(x)+\langle p,\cdot-x\rangle+\frac{1}{2r}|\cdot-\,x|^{2}$ and has vertex $v=x-rp$. Moreover, if $x\in\mathrm{Diff}^{1}(u)$, then $p=Du(x)$ is unique. This motivates the following definition.
@@
$\mathcal{C}^{+}(u,X,\tfrac{1}{r}I)\cap\mathrm{Diff}^{1}(u)$ $\xrightarrow{\,V_{r}\,}\ {\mathbb{R}}^{n}$
$x$ $\longmapsto\ I-rDu(x)$
@@
Note that the domain of $V_{r}$ consists of those points $x\in X$ which admit a global upper contact quadratic function of radius $r$ and for which $u$ is differentiable. If $u$ is (locally semi-)convex, then the upper vertex map is well-defined on all of $\mathcal{C}^{+}(u,X,\tfrac{1}{r}I)$ since $\mathcal{C}^{+}(u,X,\frac{1}{r}I)\subset\mathrm{Diff}^{1}(u)$ by Lemma 2.13 (locally semiconvex functions are differentiable at all upper contact points).
@@
For $j=1,2$, given $x_{j}\in\mathcal{C}^{+}(u,X,\frac{1}{r}I)$, let $\varphi_{j}$ be the upper contact quadratic function of radius $r$ for $u$ at $x_{j}$ with vertex $v_{j}\vcentcolon=V_{r}(x_{j})$. With the notation in Lemma 3.13, since $u$ is convex and its graph lies below those of the $\varphi_{j}$’s, then $\mathfrak{C}\subset\operatorname{epi}_{S}(u)$. Recall that $\operatorname{epi}_{S}$ is the strict epigraph, so that $z_{j}\vcentcolon=(x_{j},\varphi_{j}(x_{j}))=(x_{j},u(x_{j}))\in\Gamma(\varphi_{j})\setminus\operatorname{epi}(u)$; hence $z_{j}\in\Gamma(\varphi_{j})\setminus\mathfrak{C}$. Therefore, by Lemma 3.13, $z_{j}\notin\mathcal{H}_{j}$. It follows that the $z_{j}$’s lie on opposite sides of $\mathfrak{S}$, therefore $|x_{1}-x_{2}|\geqslant|v_{1}-v_{2}|$. ∎
@@
$K^{\delta}\vcentcolon=\{y\in K:\ d(y,\partial K)>\delta\},\quad\text{with}\ \delta\vcentcolon=\sqrt{2r\operatorname*{osc}_{K}u},$
@@
is nonempty and compact. Now, since $u$ is bounded below, its oscillation is finite. Suppose that $\delta$ is small enough to ensure that $K^{\delta}$ is not empty and fix $v\in K^{\delta}$. Clearly, if $x\in\mathcal{C}^{+}(u,K,\frac{1}{r}I,v)$, then